A copper wire has a diameter of 1.02mm,cross sectional area 8.2×10^-7 and resistivity 1.72×10^-8ohm meter.It caries a current 1.67A.Find the electric field magnitude in the wire and the potential difference between two points in the wire 50m apart.
1
Expert's answer
2020-09-30T12:28:16-0400

Electric field:

"E=\frac{\rho I}{S}=\frac{1.72×10^{-8}*1.67}{8.2×10^{-7}}=35 \frac{mV}{m}"


Potential difference:

"U=\frac{\rho*L}{S}I=\frac{1.72×10^{-8}*50*1.67}{8.2*10^{-7}}=1.75" V